In the folowing equations given write down suitable realtionship between the varaible P and Q in the form y=mx+c [striaght line graph] and identify y,x,m,c

[1] q2=ap+bp^3

[2] Q= a(p^2+b^2)/p

[3] 1/p+1/q=a

[4] p= a + q
--------
bQ
1
Expert's answer
2018-10-05T14:40:08-0400
In the following equations given write down suitable relationship between the variable P and Q in the form y=mx+c [straight line graph] and identify y,x,m,c
[1] q2=ap+bp^3
[2] Q= a(p^2+b^2)/p
[3] 1/p+1/q=a
[4] p= a + q
In this case:
c=a,m=1.